Circular No. 7564

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 A2 (LINEAR) An apparently asteroidal object discovered by LINEAR on Jan. 15 and

posted on the NEO Confirmation Page has been found to have cometary appearance on CCD images taken by P. Pravec and L. Šarounová (Ondřejov; 0′.3 coma on Jan. 16.0 UT) and by M. Tichý and M. Kočer (Kleť; diffuse, 10′′ coma on Jan. 16.9; m_{1} = 17.2). The object also links with one observed on Jan. 3 and 5.

2001 UT  α_{2000}  δ_{2000}  m_{2}
Jan. 3.31447  8^{h}39^{m}08.^{s}53  +14^{o}20′48′′.8  19.0

The astrometric observations, together with the parabolic orbital elements given below and an ephemeris, appear on MPEC 2001-B05.

T = 2001 May 24.654 TT  ω = 295.^{o}293
Ω = 295.145 }2000.0
q = 0.77977 AU  i = 36.576

SUPERNOVA 2001J IN UGC 4729 S. Beckmann and W. D. Li, University of California at Berkeley, on

behalf of LOTOSS (cf. IAUC 7514), report the discovery of an apparent supernova (mag ∼ 17.5) on unfiltered images taken on Jan. 15.4 and 16.3 UT with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 2001J is located at α = 9^{h}00^{m}43^{s}.23, δ = +17^{o}37′15′′.6 (equinox 2000.0), which is 19′′.2 east and 2′′.4 south of the nucleus of UGC 4729. A KAIT image of the same field taken on 2000 Dec. 27.4 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2001G IN MCG +08-17-43 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that spectra of SN 2001G (cf. IAUC 7560), obtained by M. Calkins on Jan. 15.31 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), show it to be a type-Ia supernova before maximum light. The spectrum consists of a blue continuum with well-defined Si II and Ca II features. The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 2 ± 2 days before maximum light.

2001 January 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green